Determine whether each of these sets is the power set of a set, where a and b are distinct elements.



1. {0}



2. {0, { a } , { 0, a }}



3. {0, { a } }



4. {0, { a }, { b }, { a, b }}

1
Expert's answer
2022-03-23T05:14:23-0400

"1:\\\left\{ \emptyset \right\} =2^{\emptyset}\,\,-\,\,yes\\2:\\\left\{ \emptyset ,\left\{ a \right\} ,\left\{ \emptyset ,a \right\} \right\} -no, the\,\,power\,\,set\,\,contains\,\,2^n\,\,elements\\3:\\\left\{ \emptyset ,\left\{ a \right\} \right\} =2^{\left\{ a \right\}}-yes\\4:\\\left\{ \emptyset ,\left\{ a \right\} ,\left\{ b \right\} ,\left\{ a,b \right\} \right\} =2^{\left\{ a,b \right\}}-yes"